# Magnetic Polarisability of Octet Baryons via Lattice QCD

### v2

@@ VIII.2 Implementation
In the process of addressing finite-volume corrections and electro quenching, we proceed as follows. We first leverage the leading loop integrals to inform the size of finite volume effects. The finite volume of the lattice requires infinite-volume integrals over momenta such as those in Eq. (81) and Eq. (82) to be calculated as sums. As such, the finite volume effects associated with the effective field theory are given by the difference between the continuous infinite-volume integrals and the discrete sums. Hence, the finite-volume corrected (FVC) valence-valence lattice polarisability is
@@
$+$ $\sum_{MB^{\prime}}\Big{(}\beta^{MB^{\prime}}_{\rm integ}(m_{\pi}^{2},\Lambda^{FV})-\beta^{MB^{\prime}}_{\rm sum}(m_{\pi}^{2},\Lambda^{FV})\Big{)}.$ ((90))
@@
where the subscripts ${v-v}$ indicate the use of the valence-valence chiral coefficients in place of the full QCD coefficients $\chi_{MB}$ and $\chi_{MB^{\prime}}$ in Eqs. Eq. (81) and Eq. (82).
@@
With the values for $a_{0}^{\Lambda}$ and $a_{2}^{\Lambda}$ determined in the fit of Eq. (91), the electro-quenching correction is incorporated by simply replacing the valence-valence coefficients by the full-QCD coefficients, i.e. by simply evaluating the full expression for the magnetic polarisability in Eq. (80). The resulting values for the octet-baryon magnetic polarisabilities at the physical pion mass $m^{\rm phys}_{\pi}=140\,$MeV are given in Tab. 7.
Systematic uncertainties associated with the chiral extrapolation are estimated through variation of the regulator parameter. The parameter is varied over the broad range $0.6\,\text{GeV}\leq\Lambda\leq 1.0\,$GeV to allow the estimation of the uncertainty associated with the higher terms of the chiral expansion. This systematic uncertainty is also included in Tab. 7.
Table 7: Magnetic polarisability values for the outer octet baryons in full QCD at the physical pion mass, $m^{\rm phys}_{\pi}=0.140\,$GeV, in infinite volume. The values include both finite-volume and electro-quenching corrections as described in the text. All values are in units of ×$\times 10^{-4}\,$fm3. Uncertainties include statistical simulation uncertainties and systematic uncertainties associated with the chiral extrapolation. These are combined in quadrature in the final column.
@@
(a) Proton extrapolation.
(b) Neutron extrapolation.
(c) $\Sigma^{+}\,$extrapolation.
(d) $\Sigma^{-}\,$extrapolation.
(e) $\Xi^{0}\,$extrapolation.
(f) $\Xi^{-}\,$extrapolation.
@@
Figure 10: Chiral extrapolation of the finite-volume- and electro-quenched-corrected (FV & EQ Corr.) octet-baryon magnetic polarisabilities to the physical pion mass indicated by the vertical dashed line. Uncorrected lattice points (Lattice Points) are horizontally offset from the corrected points. Extrapolation curves for future larger finite volumes (FV: $\cdots$ fm) are shown to illustrate the requirements for observing chiral curvature in full QCD. The curves are shown for $m_{\pi}\,L>$ 3, where $L$ is the spatial lattice length. Finally, the infinite-volume extrapolation (Inf. Vol.) relevant to nature is illustrated.
@@
It’s interesting to note that one can also use Eq. VIII.2 to explore the chiral curvature to be seen in future lattice QCD calculations. This time, one subtracts the infinite-volume integral, and adds a finite volume sum appropriate to the future lattice simulation volume. This idea is explored as we present the chiral extrapolations.
@@
Extrapolations of the octet-baryon magnetic polarisabilities to the physical point are presented in Fig. 10. The $p$, $n$, and $\Sigma^{+}$ baryons have very similar profiles as predicted by the simple quark model of Sec. II. Recall the $\Xi^{0}$ baryon is unique due to its small octet-decuplet mass splitting and containing only one $u$ quark which contributes well to the transition term of the model while minimising the negative effect in the charge distribution term by only having one $u$ quark. Together, these properties generate a large magnetic polarisability for the $\Xi^{0}$ and this is seen in our lattice QCD results.
@@
The negatively charged baryons are very interesting. We observe that they have very small magnetic polarisability values as predicted by the quark model. However with the extrapolation incorporating divergent chiral physics and including both electro-quenching corrections and finite-volume corrections, both baryons are predicted to have positive values in the infinite-volume world of full QCD. Carefully examining the results in Tab. 7, we see there is a clear prediction for the $\Sigma^{-}$ to have a positive non-vanishing magnetic polarisability. On the other hand, the positive polarisability for the $\Xi^{-}$ is only a $1\,\sigma$ effect.
@@
Figure 11: The valence-valence finite-volume (FV), electro-quenching (EQ), and total (FV+EQ) corrections for each outer octet baryon as a function of pion mass. These corrections are added to the original results from the lattice.
@@
Also of interest is the very small difference between the values of the proton and neutron. Based on our discussion of the quark model in Sec. II, we would expect the neutron to have a larger magnetic polarisability. The additional up quark in the proton acts to increase the magnitude of the negative charge distribution term. Otherwise, the transition term is identical for the proton and neutron. We will examine this difference more carefully in Sec. X.
@@
We now turn our attention to the subtle electro-quenching and finite-volume corrections for the baryons. The details of these corrections are shown in Fig. 11. With the exception of the valence-valence volume correction for the $\Sigma^{-}\,$, these corrections act to increase the magnetic polarisability. The $\Sigma^{-}\,$is unique and is discussed further below.
@@
In all cases, the dominant correction at heavy pion masses is the electro-quenching correction. However, this correction is still small. As the pion mass approaches the physical value, both volume and electro-quenching corrections increase in magnitude, but the volume corrections increase much more quickly, resulting in corrections of approximately equal magnitude at the lightest pion mass considered in this work.
@@
The $\Sigma^{-}\,$is interesting as it has large electro-quenching corrections. In the valence-valence sector, the sign of the chiral curvature is negative, such that the finite-volume correction of the valence-valence sector acts to decrease the magnetic polarisability. This is shown in the upper plot of Fig. 12. Upon implementing the electro-quenching correction, the sign of the chiral curvature becomes positive as illustrated in the lower plot of Fig. 12.
@@
This change in sign underscores the importance of the order in which the corrections are implemented. Finite-volume corrections in the valence-valence sector are implemented first as this is the frame in which the finite-volume lattice results are obtained. Once volume corrected the lattice results can be fit to determine the coefficients of the residual series. Provided one uses a physical regulator parameter, the residual series coefficients are insensitive to sea-quark-loop contributions and one can model the electro-quenching corrections by changing the coefficients of the loop-integral coefficients. With the full QCD result in hand, one can plot the finite-volume- and electro-quenched-corrected points and draw the infinite-volume extrapolation curve. From here one can also explore other finite volumes as done in Fig. 10 and Fig. 12.
@@
Figure 12: The chiral extrapolation of the finite-volume corrected valence-valence lattice QCD data (top) compared to the finite-volume and electro-quenching corrected extrapolation (bottom) for the $\Sigma^{-}\,$. The magnitude of the electro-quenching correction results in a sign change for the nonanalytic chiral curvature, a behaviour not observed for any other baryon.
@@ IX Comparison to other predictions
The Particle Data Group Workman and Others (2022) quotes values for the magnetic polarisability of the proton and neutron. The values
@@
are aggregated from a number of Compton scattering experiments. The quark model prediction that the neutron should have larger polarisability than the neutron is observed in the central values, though the large uncertainty of the neutron measurement precludes a definitive statement. Our final lattice QCD values reported in Tab. 7 compare favourably with the PDG values. Figure 13 shows the our chiral extrapolation in comparison to the PDG values and individual experiments for the proton and neutron.
Figure 13: The chiral extrapolation of volume- and electro-quenched-corrected lattice QCD results is compared with the PDG values Workman and Others (2022) and individual experiments Pasquini et al. (2019); MacGibbon et al. (1995); Blanpied et al. (2001); McGovern et al. (2013); de León et al. (2001); Beane et al. (2003); Kossert et al. (2002, 2003); Myers et al. (2014); Griesshammer et al. (2012) for the proton and neutron.
@@
Figure 14: The chiral extrapolation of volume- and electro-quenched-corrected lattice QCD results are compared with other phenomenological estimates Aleksejevs and Barkanova (2011); Gobbi et al. (1996); Tanushi et al. (2001); Scoccola et al. (1996); Deshmukh and Tiburzi (2018) for the hyperons.
@@
It is also interesting to place our results in the context of previous lattice QCD studies. Here we discuss the results of three works Lee et al. (2006); Chang et al. (2015); Bignell et al. (2020). We first note that Refs. Lee et al. (2006) and Chang et al. (2015) were carried out in the quenched approximation. In addition, Refs. Lee et al. (2006) and Chang et al. (2015) did not use a background-field corrected clover action to remove the additive mass renormalisation associated with the Wilson term when a background field is applied Bignell et al. (2019).
@@
In the early work of Ref. Lee et al. (2006) the magnetic polarisabilities of all octet baryons were determined at a range of relatively heavy pion masses. In that work, the Landau term was not considered. However, we find the energy shift associated with the Landau term to be of similar magnitude to the total energy shift and therefore we do not compare with their results for charged baryons. Accounting for their different definition of the magnetic polarisability, we divide their results by $4\pi$ and compare the $\Xi^{0}\,$and neutron magnetic polarisabilities. We find our results for $\Xi^{0}\,$to be much larger than that reported in Ref. Lee et al. (2006), approaching a factor of 3. Our results for the neutron at the lightest quark mass considered in Ref. Lee et al. (2006) are 60% larger.
@@
The observed discrepancy may be associated with the boundary conditions explored in Ref. Lee et al. (2006). While a fixed spatial boundary condition avoids the uniform magnetic field quantisation condition of Eq. (45), it is difficult to avoid the spatial boundary in the lattice QCD simulations introducing new systematic errors.
@@
The lattice simulation of Ref. Chang et al. (2015) determined the polarisability of the proton and neutron at $m_{\pi}\sim 806\,$MeV. Again, the neutron results presented there are small at approximately half that presented here. In the case of the proton, they do attempt to fit the Landau term, using their smaller field strength correlation functions to identify which Landau level the particle rests in. They fit for the $n$th Landau level, allowing $n$ to take a positive real value. As $n$ should be an integer, they round the resulting value. However, due to the magnitude of the Landau term relative to the energy shift, their approach introduces a large uncertainty into the energy shift. As such, we do not consider their proton results further. Perhaps it is worthy to note, our use of Landau mode projection at the baryon sink removes all uncertainty surrounding the Landau levels.
@@
Finally, Bignell et al. Bignell et al. (2020) determined the proton and neutron magnetic polarisabilities on the PACS-CS ensembles using similar methods which we have extended herein. The most important extensions include the single-state isolation analysis in all correlators, the consideration of several fit windows through weighted averaging, and much higher statistics. Our neutron results have a larger slope with respect to $m_{\pi}^{2}$ but extrapolate to the same value at the physical point. Increased statistics and associated methods have resolved smaller polarisability values for the proton at the two lightest quark masses considered, increasing the slope with respect to $m_{\pi}^{2}$ and producing smaller proton magnetic polarisabilities, bringing the previous value of $2.79(22)(16)\times 10^{-4}\,$fm3 to $2.12(17)(16)\times 10^{-4}\,$fm3, just outside of $1\sigma$ agreement.
@@
Looking more broadly, there are several phenomenological studies of some of the hyperons. In the absence of experimental measurements, we compare our results with theoretical models in Fig. 14. It is here that one can observe the value of our lattice QCD simulation results for the hyperons. Our uncertainties are very small on the scale of variation in model predictions.
X Proton-Neutron Magnetic Polarisability Difference

### v3

@@ VIII.2 Implementation
In the process of addressing finite-volume corrections and electroquenching, we proceed as follows. We first leverage the leading loop integrals to inform the size of finite volume effects. The finite volume of the lattice requires infinite-volume integrals over momenta such as those in Eqs. (82) and (83) to be calculated as sums. As such, the finite-volume effects associated with the effective field theory are given by the difference between the continuous infinite-volume integrals and the discrete sums. Hence, the finite-volume-corrected (FVC) valence-valence lattice polarizability is
@@
$+$ $\sum_{MB^{\prime}}\Big{(}\beta^{MB^{\prime}}_{\rm integ}(m_{\pi}^{2},\Lambda^{FV})-\beta^{MB^{\prime}}_{\rm sum}(m_{\pi}^{2},\Lambda^{FV})\Big{)},$ ((91))
@@
where the subscripts ${v-v}$ indicate the use of the valence-valence chiral coefficients in place of the full QCD coefficients $\chi_{MB}$ and $\chi_{MB^{\prime}}$ in Eqs. (82) and (83).
@@
With the values for $a_{0}^{\Lambda}$ and $a_{2}^{\Lambda}$ determined in the fit of Eq. (92), the electroquenching correction is incorporated by simply replacing the valence-valence coefficients by the full-QCD coefficients, i.e. by simply evaluating the full expression for the magnetic polarizability in Eq. (81). The resulting values for the octet-baryon magnetic polarizabilities at the physical pion mass $m^{\rm phys}_{\pi}=140\,$MeV are given in Table 7.
Systematic uncertainties associated with the chiral extrapolation are estimated through variation of the regulator parameter. The parameter is varied over the broad range $0.6\leq\Lambda\leq 1.0\,$ GeV to allow the estimation of the uncertainty associated with the higher terms of the chiral expansion. This systematic uncertainty is also included in Table 7.
Table 7: Magnetic polarizability values for the outer octet baryons in full QCD at the physical pion mass, $m^{\rm phys}_{\pi}=0.140\,$GeV, in infinite volume. The values include both finite-volume and electroquenching corrections as described in the text. All values are in units of ×$\times 10^{-4}\,$fm3. Uncertainties include statistical simulation uncertainties and systematic uncertainties associated with the chiral extrapolation. These are combined in quadrature in the final column.
@@
Figure 10: Chiral extrapolation of the finite-volume- and electroquenched-corrected (FV & EQ Corr.) octet-baryon magnetic polarizabilities to the physical pion mass indicated by the vertical dashed line. Uncorrected lattice points (Lattice Points) are horizontally offset from the corrected points. Extrapolation curves for future larger finite volumes (FV: $\cdots$ fm) are shown to illustrate the requirements for observing chiral curvature in full QCD. The curves are shown for $m_{\pi}\,L>$ 3, where $L$ is the spatial lattice length. Finally, the infinite-volume extrapolation (Inf. Vol.) relevant to nature is illustrated. (a) Proton, (b) neutron, (c) $\Sigma^{+}\,$, (d) $\Sigma^{-}\,$, (e) $\Xi^{0}\,$, and (f) $\Xi^{-}\,$extrapolation.
@@
It is interesting to note that one can also use Eq. VIII.2 to explore the chiral curvature to be seen in future lattice QCD calculations. This time, one subtracts the infinite-volume integral and adds a finite-volume sum appropriate to the future lattice simulation volume. This idea is explored as we present the chiral extrapolations.
@@
Extrapolations of the octet-baryon magnetic polarizabilities to the physical point are presented in Fig. 10. The $p$, $n$, and $\Sigma^{+}$ baryons have very similar profiles as predicted by the simple quark model of Sec. II. Recall the $\Xi^{0}$ baryon is unique due to its small octet-decuplet mass splitting and containing only one $u$ quark which contributes well to the transition term of the model while minimising the negative effect in the charge distribution term by only having one $u$ quark. Together, these properties generate a large magnetic polarizability for the $\Xi^{0}$ and this is seen in our lattice QCD results.
@@
The negatively charged baryons are very interesting. We observe that they have very small magnetic polarizability values as predicted by the quark model. However, with the extrapolation incorporating divergent chiral physics and including both electroquenching corrections and finite-volume corrections, both baryons are predicted to have positive values in the infinite-volume world of full QCD. Carefully examining the results in Table 7, we see there is a clear prediction for the $\Sigma^{-}$ to have a positive nonvanishing magnetic polarizability. On the other hand, the positive polarizability for the $\Xi^{-}$ is only a $1\,\sigma$ effect.
@@
Figure 11: The valence-valence finite-volume (FV), electroquenching (EQ), and total (FV+EQ) corrections for each outer octet baryon as a function of pion mass. These corrections are added to the original results from the lattice.
@@
Also of interest is the very small difference between the values of the proton and neutron. Based on our discussion of the quark model in Sec. II, we would expect the neutron to have a larger magnetic polarizability. The additional up quark in the proton acts to increase the magnitude of the negative charge distribution term. Otherwise, the transition term is identical for the proton and neutron. We will examine this difference more carefully in Sec. X.
@@
We now turn our attention to the subtle electroquenching and finite-volume corrections for the baryons. The details of these corrections are shown in Fig. 11. With the exception of the valence-valence volume correction for the $\Sigma^{-}\,$, these corrections act to increase the magnetic polarizability. The $\Sigma^{-}\,$is unique and is discussed further below.
@@
In all cases, the dominant correction at heavy pion masses is the electroquenching correction. However, this correction is still small. As the pion mass approaches the physical value, both volume and electroquenching corrections increase in magnitude, but the volume corrections increase much more quickly, resulting in corrections of approximately equal magnitude at the lightest pion mass considered in this work.
@@
The $\Sigma^{-}\,$is interesting as it has large electroquenching corrections. In the valence-valence sector, the sign of the chiral curvature is negative, such that the finite-volume correction of the valence-valence sector acts to decrease the magnetic polarizability. This is shown in the upper plot of Fig. 12. Upon implementing the electroquenching correction, the sign of the chiral curvature becomes positive as illustrated in the lower plot of Fig. 12.
@@
This change in sign underscores the importance of the order in which the corrections are implemented. Finite-volume corrections in the valence-valence sector are implemented first as this is the frame in which the finite-volume lattice results are obtained. Once volume corrected the lattice results can be fit to determine the coefficients of the residual series. Provided one uses a physical regulator parameter, the residual series coefficients are insensitive to sea-quark-loop contributions and one can model the electroquenching corrections by changing the coefficients of the loop-integral coefficients. With the full QCD result in hand, one can plot the finite-volume- and electroquenched-corrected points and draw the infinite volume extrapolation curve. From here one can also explore other finite-volumes as done in Figs. 10 and 12.
@@
Figure 12: The chiral extrapolation of the finite-volume-corrected valence-valence lattice QCD data (top) compared to the finite-volume and electroquenching-corrected extrapolation (bottom) for the $\Sigma^{-}\,$. The magnitude of the electroquenching correction results in a sign change for the nonanalytic chiral curvature, a behavior not observed for any other baryon.
@@ IX Comparison to other predictions
The Particle Data Group Workman and Others (2022) quotes values for the magnetic polarizability of the proton and neutron. The values
@@
are aggregated from a number of Compton scattering experiments. The quark model prediction that the neutron should have larger polarizability than the neutron is observed in the central values, though the large uncertainty of the neutron measurement precludes a definitive statement. Our final lattice QCD values reported in Table 7 compare favorably with the PDG values. Figure 13 shows our chiral extrapolation in comparison to experimental measurements and theory-based analyses for the proton Workman and Others (2022); Pasquini et al. (2019); MacGibbon et al. (1995); Blanpied et al. (2001); McGovern et al. (2013); de León et al. (2001); Beane et al. (2003); Griesshammer et al. (2016) and neutron Workman and Others (2022); Kossert et al. (2002, 2003); Myers et al. (2014); Griesshammer et al. (2012, 2016).
Figure 13: The chiral extrapolation of volume- and electroquenched-corrected lattice QCD results for the proton (top) and neutron (bottom) are compared with PDG averages Workman and Others (2022), experimental results and theory-based analyses. Results at the physical point are offset for clarity. Blanpied et al. Blanpied et al. (2001), MacGibbon et al. MacGibbon et al. (1995), Olmos de Leon et al. de León et al. (2001) and Kossert et al. Kossert et al. (2003) report experimental results. Myers et al. Myers et al. (2014) provide a result from deuteron scattering with effective field theory input. McGovern et al. McGovern et al. (2013), Beane et al. Beane et al. (2003) and Griesshammer et al. Griesshammer et al. (2016) report effective-field-theory based analyses. Pasquini et al. Pasquini et al. (2019) utilizes dispersion relations to obtain the result presented.
@@
Figure 14: The chiral extrapolation of volume- and electroquenched-corrected lattice QCD results are compared with other phenomenological estimates Aleksejevs and Barkanova (2011); Gobbi et al. (1996); Tanushi et al. (2001); Scoccola et al. (1996); Deshmukh and Tiburzi (2018) for the hyperons.
@@
It is also interesting to place our results in the context of previous lattice QCD studies. Here we discuss the results of three works Lee et al. (2006); Chang et al. (2015); Bignell et al. (2020). We first note that Refs. Lee et al. (2006) and Chang et al. (2015) were carried out in the quenched approximation. In addition, Refs. Lee et al. (2006) and Chang et al. (2015) did not use a background-field-corrected clover action to remove the additive mass renormalization associated with the Wilson term when a background field is applied Bignell et al. (2019).
@@
In the early work of Ref. Lee et al. (2006) the magnetic polarizabilities of all octet baryons were determined at a range of relatively heavy pion masses. In that work, the Landau term was not considered. However, we find the energy shift associated with the Landau term to be of similar magnitude to the total energy shift, therefore we do not compare with their results for charged baryons. Accounting for their different definition of the magnetic polarizability, we divide their results by $4\pi$ and compare the $\Xi^{0}\,$and neutron magnetic polarizabilities. We find our results for $\Xi^{0}\,$to be much larger than that reported in Ref. Lee et al. (2006), approaching a factor of 3. Our results for the neutron at the lightest quark mass considered in Ref. Lee et al. (2006) are 60% larger.
@@
The observed discrepancy may be associated with the boundary conditions explored in Ref. Lee et al. (2006). While a fixed spatial boundary condition avoids the uniform magnetic-field quantization condition of Eq. (45), it is difficult to avoid the spatial boundary in the lattice QCD simulations introducing new systematic errors.
@@
The lattice simulation of Ref. Chang et al. (2015) determined the polarizability of the proton and neutron at $m_{\pi}\sim 806\,$MeV. Again, the neutron results presented there are small at approximately half that presented here. In the case of the proton, they do attempt to fit the Landau term, using their smaller field strength correlation functions to identify which Landau level the particle rests in. They fit for the $n$th Landau level, allowing $n$ to take a positive real value. As $n$ should be an integer, they round the resulting value. However, due to the magnitude of the Landau term relative to the energy shift, their approach introduces a large uncertainty into the energy shift. As such, we do not consider their proton results further. Perhaps it is worthy to note, our use of Landau-mode projection at the baryon sink removes all uncertainty surrounding the Landau levels.
@@
Finally, Bignell et al. Bignell et al. (2020) determined the proton and neutron magnetic polarizabilities on the PACS-CS ensembles using similar methods which we have extended herein. The most important extensions include the single-state isolation analysis in all correlators, the consideration of several fit windows through weighted averaging, and much higher statistics.
It is the latter consideration that has led to a notable change in the magnetic polarizability of the proton. The current work has 8 times the statistics and this has enabled the explicit examination of the excited-state contamination in each of the correlation functions prior to combining them to obtain the polarizability. This contrasts the analysis possible in Bignell et al. Bignell et al. (2020) where excited-state contamination was examined only in the ratio after combining the correlation functions to access the magnetic polarizability energy shift $\delta E_{\beta}$.
We have now discovered the proton’s correlation function ratio displays an early plateau in $\delta E_{\beta}$ before the underlying correlation functions have reached single-state isolation. Thus, the results in Ref. Bignell et al. (2020) are based on fit windows commencing at earlier Euclidean times.
For example at $m_{\pi}^{\rm PACS-CS}=296\,$MeV, we showed in Sec. VI and Fig. 4 that the underlying correlation functions for the proton do not reach single-state isolation until $t=24$. In Ref. Bignell et al. (2020), the fit window $[t_{s},t_{f}]=[20,24]$ was utilized. Excited-state contamination gives rise to a difference in the results reported.
This behavior is shared by the neutron, though to a lesser degree as $\delta E_{\beta}$ for the neutron tends to show plateaulike behavior only later in Euclidean time, allowing the underlying correlation functions to have settled further to single-state isolation. Still, the neutron results presented here have a larger slope with respect to $m_{\pi}^{2}$ but extrapolate to a similar value at the physical point.
Increased statistics and weighted averaging of acceptable fit windows have resolved smaller polarizability values for the proton at the two lightest quark masses considered, increasing the slope with respect to $m_{\pi}^{2}$ and producing smaller proton magnetic polarizabilities at the physical point. These systematic advances bring the previous value of $2.79(22)(16)\times 10^{-4}$ to $2.12(17)(16)\times 10^{-4}\,$fm3, just outside $1\sigma$ agreement.
@@
Looking more broadly, there are several phenomenological studies of the hyperons. In the absence of experimental measurements, we compare our results with theoretical models in Fig. 14. It is here that one can observe the value of our lattice QCD simulation results for the hyperons. Our uncertainties are very small on the scale of variation in model predictions.
X Proton-Neutron Magnetic Polarizability Difference